find all real numbers excluded from the domain of
{ { { f(x)= x-6/5x^2+7x-6 } } }
the give the domain using set notation

Answer by venugopalramana(3286)

SEE THE FOLLOWING EXAMPLE AND TRY..IF STILL IN DIFFICULTY PLEASE COME BACK
Find all real numbers excluded from the domain of
{ { { f(x)= 16x^2-9/20x^2-23x+6 } } }
then give the domain using set notation.
f(x)= 16x^2-9/20x^2-23x+6
={(4X)^2-(3)^2}/{20X^2-8X-15X+6}
=(4X-3)^2/{4X(5X-2)-3(5X-2)}
=(4X-3)^2/(4X-3)(5X-2)
HENCE THE DOMAIN IS ALL REAL NUMBERS EXCEPT 4X-3=0..OR...4X=3...OR..X=3/4 AND
5X-2=0...OR....5X=2...OR....X=2/5.......
HENCE DOMAIN OF F(X)={X E R (X IS NOT EQUAL TO 3/4 AND 2/5)}